TGBT14A _912_116


Joe and Adam are testing the most up to date BOF files, which should have almost all fixes in place (bi modal behaviour, shifted center frequency, better/flatter internal filters, etc...). Tests are all in L-band

L8/LBW1 v191 (modes 10-19)

Data Guide

  • Scan 01 - Mode11_tp_Lband - 60s - Tsys 106K
Fix for the Balance not implemented, rerunning scan
  • Scan 02 - Mode11_tp_Lband - 60s - Tsys 107K
  • Scan 03 - Mode16_tp_Lband - 60s - (to see if the Tsys is BOF related) - Not a good BOF to test in current Balance fix situation
  • Scan 04 - Mode09_tp_Lband - 60s - high tsys - Noticed the boom is out
  • Scan 05 - Mode09_tp_Lband - 60s -
  • Scan 06 - Mode11_tp_Lband - 60s -
  • Scan 07 - Mode12_fsw_Lband - 60s - fsw to remove baselines and look for linearity All looks good!
  • Scan 08 through 12 - Mode12_fsw_Lband - 60s - a search for bi-modality and general stability
  • Scan 13 - Mode24_fsw_Lband - 60s - all data is very close, though there is a slight variation in the FS data in pairs of two, leading to very small changes in the spectra and system temperatures measured. I believe this is due to the way the system is balanced in groups of two subwindows at a time for both polarizations.
Topic revision: r1 - 2015-08-12, AdamKobelski
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback